Help me out

(1) Solve the right triangle where A = 40°, b = 6
Answer #1

Let us assume the triangle is \Delta ABC , such that 06 => and <A=40^0 , with a,b,c being the sides opposite angles A, B and C respectively

We know, in a triangle, the angles add up to 180 degrees

<A+<B+<C=180^0

40^0+90^0+<C=180^0

<C=180^0-130^0

<C=50^0

Also, the question specifies, b=6

As per the law of sines,

a/SinA=b/SinB=c/SinC

or

a/SinA=b/SinB

a/Sin40^0=6/Sin90^0

a=6*Sin40^0/Sin90^0

a=6*0.64278760968 /1

a=6*0.64278760968 /1=3.856

b/SinB=c/SinC

6/Sin90^0=c/Sin50^0

c=6*Sin50^0/Sin90^0

c=6*0.76604444311/1

c=4.596

Hence

<A=40^0,<B=90^0,<C=50^0,a=3.856,b=6,c=4.596